OpenStudy (anonymous):

URGENT! Please show work! A teacher is training for a half marathon in 2013. In one week she ran 5km on the first day and increased her distance by 10% on each subsequent day. a) How far did she run on the seventh day? b) What is the total distance she ran during the week?

OpenStudy (anonymous):

Do you know compound interest?

OpenStudy (anonymous):

I might need a refresher :(

OpenStudy (anonymous):

the way i first tried solving this was with a geometric series formula. unless its different, im plugging everything in wrong...

OpenStudy (anonymous):

currently i am using this formula a of n=a of 1 times (r)^n-1

OpenStudy (anonymous):

\[\sum_{n=0}^{7} 5*(1.1)^{n} = 5*(1.1)^{0} + 5*(1.1)^{1} +5*(1.1)^{2}....\]

OpenStudy (anonymous):

n=6 is the seventh day in this set up, little confusing there sorry.

OpenStudy (anonymous):

thats all right, the main thing im confused about is what formula I use to solve it?

OpenStudy (anonymous):

i have specific term, finite sum, or infinite

OpenStudy (anonymous):

and the total for the week is not \[\sum_{n=0}^{7} \] but rather \[\sum_{n=0}^{6}\]

OpenStudy (anonymous):

alright, let me try that out

OpenStudy (anonymous):

the answer to the problem is 8.86 for a and 47.44 for b btw

OpenStudy (anonymous):

\[5*(1.1)^{6} = ?\]

OpenStudy (anonymous):

8.86 :D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks but if i dont mind asking how did you get 1.1 again?

OpenStudy (anonymous):

i think i understand everything else

OpenStudy (anonymous):

use this for the sum \[a*(\frac{ 1-r^6 }{ 1-r })\]

OpenStudy (anonymous):

where a=5 and r=1.1

OpenStudy (anonymous):

oh alrighty

OpenStudy (anonymous):

thank you so much, i really appreciate all the help!

OpenStudy (anonymous):

sure:)